calculate the percentage of available chlorine in a sample of 3.546 g of bleaching powder which was dissolved in 100 mL of water,25 mL of this solution,on treatment with KI and dilute acid,required 20 mL of 0.125 N sodium thiosulphate solution.

Grade:11

1 Answers

ashutosh
5 years ago
%Cl2 = 3.55*x*V (ml)/W (g),   x = molarity of hypo solution, V =volume of hypo solution used in titration
         = 3.55*0.125*20/3.546
         = 25 %